### v1

@@ 4 DYNAMIC ORCHESTRATION: INTER-DOMAIN SPACE DIGITAL TWIN
As different CSN architectures provide varying advantages in terms of coverage and link reliability depending on dynamic operational conditions and mission requirements, a mechanism for intelligent and dynamic switching across lunar communication scenarios is critical. Therefore, we propose a digital twin that serves as an dynamic decision-making engine, modeling the entire communication system from lunar nodes to ground stations on Earth via LLO and GEO relays.
The primary goal of the digital twin is to optimize communication strategies during missions by balancing the crucial need for reliability with effective resource utilization, particularly in terms of transmit power, which improves overall system performance and enhances adaptability. Here, the inter-domain space digital twin is updated with the status, link conditions, and operational parameters of all relevant nodes each minute. Its comprehensive awareness allows to proactively assess and optimize for continuous communication, striving for stable reliability performance, and efficiency even when faced with potential risks or dynamic channel variations. This capability underpins its function of providing automated and informed decision making for network operations.
@@ 5 NUMERICAL RESULTS
cells = c,m,
row17 = c,row18 = c,row19 = c,
column5 = c,column6 = c,column7 = c,column8 = c,
cell11 = c=2c,cell13 = c,cell14 = c,
cell21 = c=2c,cell23 = c,cell24 = c,
cell31 = c=2c,cell34 = c=6,
cell41 = r=4c,cell42 = c,cell43 = c,cell44 = c,
cell54 = c, cell64 = c, cell74 = c,
cell81 = r=7c,cell82 = c,cell83 = c,cell84 = c,
cell94 = c, cell104 = c, cell114 = c,
cell124 = c, cell134 = c, cell144 = c,
cell151 = c=2c,
cell161 = c=2,cell164 = c=6,
cell171 = c=2c,
vlines, hline1-4,8,15-18 = -, hline5-7,9-15 = 2-9,
Parameters& Unit
Moon
to
Earth
Moon
to
LLO
LLO
to
Earth
Moon
to
GEO
GEO
to
Earth
LLO
to
GEO
GHz 26.25 27.25 26.25 27.25 26.25 27.25
MHz 50 50 50 50 50 50
Transmitter
W 2 0.5 2 2 0.01 2
0.85 0.85 0.90 0.85 0.93 0.90
m 1.5 1.5 1.5 1.5 4.6 1.5
dB 1.5 1.5 1 1.5 0.5 1
Receiver
0.95 0.90 0.95 0.93 0.95 0.93
m 34 1.5 34 4.6 34 4.6
dB 0.5 1 0.5 0.5 0.5 0.5
K 50 100 50 70 50 70
K 290 250 290 270 290 270
K 290 250 290 270 290 270
0.95 0.90 0.95 0.95 0.95 0.95
K 290 290 290
K 2.725 2.725 2.725 2.725 2.725 2.725
Rician- dB 10 20 15 13 30 40
@@ 5.1 Simulation Set-up and Configuration
The Mons Malapert ()222The selenographic coordinate system is utilized to specify Mons Malapert’s location on the lunar surface., a promising candidate for future missions because of its potential water ice presence and near-consistent illumination [44], hosts the lunar communication terminal (LCT), as shown in Fig. 2.
@@
Two-way communication is utilized by all relay satellites, which are equipped with antenna that can operate in both directions.
Since the Moon lacks an atmosphere like Earth’s and the lunar regolith reflects light differently than on Earth, we conduct our analyses by examining various illumination conditions of the lunar surface. These illumination fluctuations may also impact receiver antennas pointing toward an LLO satellite, exposing them to the lunar surface in the background. Therefore, the brightness temperature for antennas pointing to the Moon or the LLO satellite varied as , while it is fixed at for antennas pointing to GEO satellites.
@@ 5.2 In-Orbit Link Geometry and Propagation Dynamics
We observe daily intervals of visibility and non-visibility for links in Fig. 3, as well as acquire insight into LoS conditions through changes in azimuth and elevation. The elevation angles of Moon-to-Earth direct links are usually around five degrees, denoting that the Moon is barely above the horizon from the perspective of three GS. Although there are still LoS, the fluctuating and low elevation means highly variable atmospheric path lengths for ground stations. When this is combined with ranges from roughly km to over km, the need for alternative relay solutions becomes clear. Moon-to-LLO links offer very short ranges, which are advantageous for minimizing path loss on the lunar access segment, but they have highly dynamic azimuth and elevation angles due to multiple LLO satellites. However, this benefit is countered by the rapid changes in azimuth and elevation, indicating a highly dynamic link and demanding competent antenna pointing and frequent handovers to maintain connectivity. LLO-to-Earth links exhibit higher azimuth and elevation compared to the Moon-to-Earth link, but at larger distances. These longer and highly variable distances, coupled with dynamic angles, mean significantly fluctuating path loss.
@@ 5.3 Architecture Reliability Assessment
This section examines the reliability performance of four lunar communication architectures with corresponding scenarios, building upon the dynamic link conditions established in Fig. 3. Fig. 4 illustrates the outage probability for each scenario over time under different brightness temperatures and an SNR threshold of 10 dB. Individual architecture analyses present their robustness for inter-domain space communication and implicitly lay the groundwork for the requirement of a dynamic decision-making mechanism among them.
@@
The outage probability for direct Moon-to-Earth (scenario 1) communication exhibits significant variability and frequently reaches high values, as shown in Fig. 4a. This variability directly results from the visibility constraints observed in Fig. 3, where the elevations are very low, despite the presence of LoS via three ground stations. The high outage probabilities, particularly during high ranges, demonstrate that conventional direct communication is not sufficiently reliable for continuous operations on future lunar missions. The impact of increasing illumination is visible; while it influences the noise floor, the primary driver of outage in this architecture is the fundamental geometric and visibility limitation.
@@
Scenario 2, utilizing LLO relays, shows improved reliability at first look in Fig. 4b due to the fewer colored points compared to direct communication. In fact, the number of high outages is larger, especially mostly around the times of handover between LLO satellites and with higher . This shows the dynamic nature of LLO links, as shown in Fig. 3, where rapid changes in visibility and frequent handovers might lead to instant changes for the links or outages. LLO relays offer lower path loss and better visibility in the cislunar domain (due to closer proximity to the lunar surface); however, the overall reliability depends on the availability and stability of link hops in the architecture. This means that maintaining a consistently low outage across both hops is challenging due to inherent orbital dynamics.
@@
Scenario 3, relying on bidirectional GEO relay, generally demonstrates a more stable and often lower outage probability compared to scenario 1 and scenario 2. The outage probabilities mostly fall at and below, indicating significantly higher reliability. The smooth and periodic link dynamics of GEO satellites observed in Fig. 3 contribute to this improved stability. While higher values lead to an increase in the outage probability, the overall reliability is maintained at lower levels than other scenarios. Both results, in Fig. 3 and Fig. 4c, justify the advantages of the proposed NSN architecture for maintaining consistent and robust connectivity. over large distances.
@@
The multi-hop scenario 4 generally exhibits the lowest outage probabilities among all scenarios across various values, except for the highest outages caused during the handover of LLO relay satellites as shown in Fig. 4d. This justifies that if this problem is overcome (such as by using more LLO satellites), the architecture will provide superior reliability by leveraging the benefits of both LLO (for short ranges and better visibility) and GEO for robust backhaul to Earth (stable and predictable link dynamics). On the other hand, boosting the multi-hop design with more satellites inherently means more redundancy, complexity, and resource usage. Therefore, the challenging trade-off between performance and cost constraints leads to scenario 3 being more amenable, while the higher performance demands make scenario 4 superior.
@@
Across all scenarios, an increase in the brightness temperature typically correlates with an increase in outage probability. This is because contributes to the external antenna noise temperature, which in turn increases the operational equivalent noise temperature and thus the total noise power at the receiver. The sensitivity of outage probability to lunar surface brightness emphasizes the importance of accurately modeling the noise environment. Our comprehensive reliability analysis, showing the varied performance of each architecture, then establishes the requirement for dynamic management to maintain stable communication. This justifies the motivation for the development of the inter-domain space digital twin, as results show that no single architecture is always ideal and that a dynamically adaptive solution is necessary to overcome the inherent trade-offs between them.
@@ 5.4 Validation of Dynamic Orchestration
We observe the dynamic and complicated nature of lunar communication with significant variability in link geometries and challenges like limited line-of-sight, rapidly changing LLO relay links, and path losses via Fig. 3. Fig. 4 further reveals no single architecture consistently provides optimal reliability. This comprehensive analysis establishes that a dynamic orchestration is indispensable for continuous, robust, and resource-efficient lunar communication. Therefore, we introduce an inter-domain space digital twin that continuously models the entire communication system, dynamically optimizing strategies by balancing unwavering reliability with efficient transmit power utilization.
Given overall reliability performance with a relatively low requirement in Fig. 4, the third scenario—based on GEO relaying communication architecture—is set as a reserved fallback scenario. The practical efficacy of this digital twin is compellingly demonstrated in Fig. 5, which illustrates the reliability performance throughout a day under varying brightness temperatures. For a brightness temperature of 200 K, as shown in Fig. 5(a), the digital twin successfully maintains the communication link predominantly within the high reliability target (). The bar, showing chosen scenarios by minutes, indicates dynamic switching, primarily between scenario 3 and scenario 4, to achieve this consistent high reliability. This is crucial for mission success but also security, as it enables real-time command and control and vital data transmission, even under dynamic lunar conditions. The capability to achieve this while optimizing power consumption is pivotal, given that power is a scarce resource in the cislunar domain. In Fig. 5(a), we also observe occasional moderate reliability cases, but the system quickly recovers, and the fallback scenario (scenario 3) is rarely engaged. This highlights that our proposed inter-domain space digital twin can adapt to dynamic conditions (as in Fig. 3) and leverage architectural strengths for superior performance.
In Fig. 5(b), under an increased brightness temperature of 400 K, the digital twin continues to strive for high reliability. While largely operating within the high reliability level, more frequent and slightly longer periods show outage probability falling into moderate () or low reliability () levels. The fallback scenario is also utilized more often compared to Fig. 5(a) ( K). This demonstrates the digital twin’s robustness in adapting to more challenging noise environments, dynamically selecting scenarios to mitigate performance degradation. This dynamism is essential to overcome traditional communication limitations and ensures unwavering reliability, paramount for mission success and safety, even when maintaining the absolute highest reliability becomes more power intensive or geometrically constrained due to scarce power resources.

### v2

@@ 4 DYNAMIC ORCHESTRATION: INTER-DOMAIN SPACE DIGITAL TWIN
As different CSN architectures provide varying advantages in terms of coverage and link reliability depending on dynamic operational conditions and mission requirements, a mechanism for intelligent and dynamic switching across lunar communication scenarios is critical. Therefore, we propose a digital twin, the dynamic decision-making engine illustrated in the right panel of Fig. 1, which models the entire communication system from lunar nodes to ground stations on Earth via LLO and GEO relays.
The primary goal of the digital twin is to optimize communication strategies during missions by balancing the crucial need for reliability with effective resource utilization, particularly in terms of transmit power, which improves overall system performance and enhances adaptability. Here, the inter-domain space digital twin is updated with the status, link conditions, and operational parameters of all relevant nodes each minute. Its comprehensive awareness allows to proactively assess and optimize for continuous communication, striving for stable reliability performance, and efficiency even when faced with potential risks or dynamic channel variations. This capability underpins its function of providing automated and informed decision-making for network operations.
@@ 5 NUMERICAL RESULTS
| Parameters | Unit | Moon | ||||||
| to | ||||||||
| Earth | Moon | |||||||
| to | ||||||||
| LLO | LLO | |||||||
| to | ||||||||
| Earth | Moon | |||||||
| to | ||||||||
| GEO | GEO | |||||||
| to | ||||||||
| Earth | LLO | |||||||
| to | ||||||||
| GEO | ||||||||
| GHz | 26.25 | 27.25 | 26.25 | 27.25 | 26.25 | 27.25 | ||
| MHz | 50 | 50 | 50 | 50 | 50 | 50 | ||
Transmitter |
W | 2 | 0.5 | 2 | 1.99 | 0.01 | 2 | |
| 0.85 | 0.85 | 0.90 | 0.85 | 0.93 | 0.90 | |||
| m | 1.5 | 1.5 | 1.5 | 1.5 | 4.6 | 1.5 | ||
| dB | 1.5 | 1.5 | 1 | 1.5 | 0.5 | 1 | ||
Receiver |
0.95 | 0.90 | 0.95 | 0.93 | 0.95 | 0.93 | ||
| m | 34 | 1.5 | 34 | 4.6 | 34 | 4.6 | ||
| dB | 0.5 | 1 | 0.5 | 0.5 | 0.5 | 0.5 | ||
| K | 50 | 100 | 50 | 70 | 50 | 70 | ||
| K | 290 | 250 | 290 | 270 | 290 | 270 | ||
| K | 290 | 250 | 290 | 270 | 290 | 270 | ||
| 0.95 | 0.90 | 0.95 | 0.95 | 0.95 | 0.95 | |||
| K | 290 | 290 | 290 | |||||
| K | 2.725 | 2.725 | 2.725 | 2.725 | 2.725 | 2.725 | ||
| Rician- | dB | 10 | 20 | 15 | 13 | 30 | 40 |
@@ 5.1 Simulation Set-up and Configuration
The Mons Malapert ()222The selenographic coordinate system is utilized to specify Mons Malapert’s location on the lunar surface., a promising candidate for future missions because of its potential water ice presence and near-consistent illumination [baris_hybrid], hosts the lunar communication terminal (LCT), as shown in Fig. 2.
@@
Two-way communication is utilized by all relay satellites, which are equipped with antennas that can operate in both directions.
Since the Moon lacks an atmosphere like Earth’s and the lunar regolith reflects light differently than on Earth, Moon’s actual brightness temperature varies spatially and temporally. These illumination fluctuations may also impact receiver antennas pointing toward an LLO satellite, exposing them to the lunar surface in the background. To encompass this range of conditions in our analysis, the brightness temperature for antennas pointing toward the Moon or the LLO satellite was varied as: , while it is fixed at for antennas pointing to GEO satellites.
The minimum SNR threshold is set to 10 dB, a value that supports high-fidelity communication for the robust modulation and coding schemes typically required for cislunar and deep space missions.
We established three targeted reliability levels: with corresponding maximum acceptable outage probability targets of , respectively.
@@ 5.2 In-Orbit Link Geometry and Propagation Dynamics
We observe daily intervals of visibility and non-visibility for links in Fig. 3, as well as acquire insight into LoS conditions through changes in azimuth and elevation. The elevation angles of Moon-to-Earth direct links are usually around five degrees, denoting that the Moon is barely above the horizon from the perspective of three GS. Although there are still LoS, the fluctuating and low elevation means highly variable atmospheric path lengths for ground stations. When this is combined with ranges from roughly km to over km, the need for alternative relay solutions becomes clear. Moon-to-LLO links offer very short ranges, which are advantageous for minimizing path loss on the lunar access segment, but they have highly dynamic azimuth and elevation angles due to multiple LLO satellites. However, this benefit is countered by the rapid changes in azimuth and elevation, indicating a highly dynamic link and demanding precise antenna pointing and frequent handovers to maintain connectivity. LLO-to-Earth links exhibit higher azimuth and elevation compared to the Moon-to-Earth link, but at larger distances. These longer and highly variable distances, coupled with dynamic angles, mean significantly fluctuating path loss.
@@ 5.3 Architecture Reliability Assessment
This section examines the reliability performance of four lunar communication architectures with corresponding scenarios, building upon the dynamic link conditions established in Fig. 3. Fig. 4 illustrates the outage probability for each scenario over time under different brightness temperatures and an SNR threshold of 10 dB. Individual architecture analyses present their robustness for inter-domain space communication and implicitly lay the groundwork for the requirement of a dynamic decision-making mechanism among them. These analyses also allow us to address the trade-off between each scenario’s architectural complexity and its resulting operational robustness.
@@
While simple in its DTE network design, scenario 1 requires immense complexity on the ground to overcome LoS issues and the strong signal attenuation. This trade-off is compounded by its poor operational robustness, as shown in Fig. 4a. The outage probability for direct Moon-to-Earth communication exhibits significant variability and frequently reaches high values. This variability directly results from the visibility constraints observed in Fig. 3, where the elevations are very low, despite the presence of LoS via three ground stations. The high outage probabilities, particularly during high ranges, demonstrate that conventional direct communication is not sufficiently reliable for continuous operations on future lunar missions. The impact of increasing illumination is visible; while it influences the noise floor, the primary driver of outage in this architecture is the fundamental geometric and visibility limitation.
@@
Scenario 2 adds architectural complexity with LLO relays to solve LoS issues in scenario 1, but it introduces new robustness challenges. In fact, the number of high outages is larger in Fig. 4b, especially mostly around the times of handover between LLO satellites and with higher . This shows the dynamic nature of LLO links, as shown in Fig. 3, where rapid changes in visibility and frequent handovers might lead to instant changes for the links or outages. LLO relays offer lower path loss and better visibility in the cislunar domain (due to closer proximity to the lunar surface); however, the overall reliability depends on the availability and stability of link hops in the architecture. This means that maintaining a consistently low outage across both hops is challenging due to inherent orbital dynamics.
@@
Scenario 3, relying on bidirectional GEO relay, generally demonstrates a more stable and often lower outage probability compared to scenario 1 and scenario 2. The outage probabilities mostly fall at and below, indicating significantly higher reliability. The smooth and periodic link dynamics of GEO satellites, as observed in Fig. 3, contribute to this improved stability. While our analysis uses complex DSN ground stations for all scenarios to ensure a fair comparison—making scenario 3 initially appear more complex—a DSN ground station is not essential for the stable GEO-to-Earth link. This highlights another potential feasibility of the architecture in practice: using simpler NSN ground stations to offload immense hardware complexity from the ground, thereby creating a more scalable solution. Although the outage probability increases with higher values, the overall reliability remains lower than in other scenarios. Both results, in Fig. 3 and Fig. 4c, justify the advantages of the proposed NSN architecture for maintaining consistent and robust connectivity over large distances.
@@
As the most complex architecture, the multi-hop scenario 4 generally exhibits the lowest outage probabilities among all scenarios across various values, except for the highest outages caused during the handover of LLO relay satellites, as shown in Fig. 4d. This justifies that if this problem is overcome (such as by using more LLO satellites), the architecture will provide superior reliability by leveraging the benefits of both LLO (for short ranges and better visibility) and GEO for robust backhaul to Earth (stable and predictable link dynamics). On the other hand, boosting the multi-hop design with more satellites inherently means more redundancy, complexity, and resource usage. Therefore, the challenging trade-off between performance and cost constraints leads to scenario 3 being more amenable, while the higher performance demands make scenario 4 superior.
@@
Across all scenarios, an increase in the brightness temperature typically correlates with an increase in outage probability. This is because contributes to the external antenna noise temperature, which in turn increases the operational equivalent noise temperature and thus the total noise power at the receiver. The sensitivity of outage probability to lunar surface brightness emphasizes the importance of accurately modeling the noise environment.
While our analysis assumes a uniform surface brightness temperature for the 24-hour period since the average temperature at a location like Mons Malapert does not fluctuate dramatically within this short time frame, the broader context of the 29.5-day lunar cycle would impose slow variations on . Furthermore, the real-world lunar thermal environment is complicated by spatial and topographical variations that create a complex thermal map. Although the LCT is not mobile, its location at Mons Malapert near the South Pole is an area known for long shadows and sharp temperature gradients. An antenna tracking a satellite could therefore sweep its view from a sunlit patch to a permanently shadowed crater, causing rapid and unpredictable fluctuations in the external antenna noise temperature. These thermal hot spots would impact the outage probability momentarily due to noise rather than link geometry alone. These real-world complexities create a critical gap that static communication architectures cannot fill.
This is precisely the motivation for the development of the inter-domain space digital twin. The results show that no single architecture is always ideal, and a dynamically adaptive solution is necessary to overcome the inherent trade-offs between them. The digital twin is designed not just to select the best path based on predictable orbital dynamics but to proactively mitigate these time- and location-dependent noise variations. By integrating real-time telemetry with predictive models, the digital twin can switch to a more robust communication path before noise compromises the primary link, ensuring the unwavering reliability that future lunar missions demand.
@@ 5.4 Validation of Dynamic Orchestration
We observe the dynamic and complicated nature of lunar communication with significant variability in link geometries and challenges like limited line-of-sight, rapidly changing LLO relay links, and path losses via Fig. 3. Fig. 4 further reveals no single architecture consistently provides optimal reliability. This comprehensive analysis establishes that a dynamic orchestration is indispensable for continuous, robust, and resource-efficient lunar communication. Therefore, we introduce an inter-domain space digital twin that continuously models the entire communication, dynamically optimizing strategies by balancing unwavering reliability with efficient transmit power utilization.
| Metrics | |||||
| Scenario 1 (Direct to Earth) | |||||
| Total Outage | 10.2284 | 12.1227 | 13.6884 | 15.0746 | 16.3272 |
| Avg. | 2 | ||||
| Scenario 2 (via an LLO relay) | |||||
| Total Outage | 37.8367 | 45.2522 | 51.6477 | 56.6411 | 61.0316 |
| Avg. | 2.5 | ||||
| Scenario 3 (via a GEO relay) | |||||
| Total Outage | 0.0102 | 0.0725 | 0.2720 | 0.7117 | 1.4501 |
| Avg. | 2 | ||||
| Scenario 4 (via an LLO and a GEO relay) | |||||
| Total Outage | 30.9685 | 49.5117 | 73.6876 | 94.4604 | 119.6084 |
| Avg. | 2.51 | ||||
| All Scenarios with Digital Twin Management | |||||
| Total Outage | 0.00013 | 0.00026 | 0.00042 | 0.00073 | 0.0013 |
| Avg. | 2.0007 | 2.0014 | 2.0032 | 2.0046 | 2.0060 |
Table 2 presents a quantitative summary of the daily total outage duration and average total transmit power () over the 24-hour period. The results clearly illustrate the significant disparity in operational robustness among the individual architectures and reveal the profound improvements achieved through the dynamic orchestration of the proposed digital twin. Compared to its counterparts, the third scenario (based on GEO relaying) demonstrates the highest reliability, with the lowest outage duration for an average of 2 W. Despite its superior performance, uninterrupted connectivity is essential for cislunar missions, where even an outage of a few seconds can have critical consequences. Here, the digital twin’s role becomes mandatory: under challenging noise conditions (), it reduces the total daily outage from scenario 3’s 87 seconds to just 0.078 seconds, a significant gain in reduction of over 99.9%, all for a negligible 0.3% increase in average total power. In other words, it increases the daily link availability from 99.90% for the best static scenario to 99.9999%. This result underscores that dynamic architectural management is essential for inter-domain space networks, allowing the system to adapt to changing conditions and always select the most reliable and power-efficient communication path. The subsequent analysis addresses precisely how the digital twin achieves the dramatic performance improvements in Table 2.
Fig. 5 compellingly illustrates this minute-by-minute orchestration, showing how the digital twin sustains a highly reliable link throughout the day under different thermal noise conditions. For a brightness temperature of 200 K, as shown in Fig. 5(a), the digital twin maintains the communication link predominantly within the high-reliability target (). The lower bar reveals the system’s underlying strategy: while it generally relies on the robust and relatively simple scenarios 1 and 3, its sophistication is evident in its tactical use of more complex multi-hop architectures. For instance, in the window between minutes 1032 and 1080, a period where scenarios 1 and 3 face a higher risk of outage, the digital twin switches to scenario 4 to ensure uninterrupted connectivity. This decision illustrates the calculated trade-off at the core of the engine’s logic: it embraces higher architectural complexity only when it is essential for maintaining robustness, while simultaneously selecting the most power-efficient path that meets the reliability target. This adaptable balancing of reliability against complexity and power consumption is crucial for enabling stable real-time command and control in the resource-scarce cislunar domain. Even when conditions cause brief dips into moderate reliability, the system quickly recovers, showcasing its ability to leverage the strengths of each architecture in response to the dynamic environment illustrated earlier in Fig. 3.
To further validate the digital twin’s resilience, Fig. 5(b) shows its performance under the more challenging condition of an increased brightness temperature of 400 K, where the engine continues to strive for high reliability. While largely operating within the high-reliability level, more frequent and slightly longer periods of outage probability fall into moderate or low-reliability levels. This behavior is a direct result of the elevated noise floor accentuating the inherent architectural trade-offs discussed previously. For instance, the immense path loss associated with the otherwise stable Moon-to-GEO link (scenario 3) becomes more critical for the digital twin’s decision. Simultaneously, the system becomes more sensitive to the highly dynamic nature of LLO links, where outages are most common during the frequent satellite handovers required for continuous connectivity (a key weakness of scenarios 2 and 4). However, the digital twin adapts to more challenging noise environments and ensures robustness by dynamically selecting scenarios to mitigate performance degradation. Crucially, even as the system is pushed into moderate or low reliability more often, it successfully avoids the fallback reserved level entirely. This dynamic adaptability allows the system to effectively manage the trade-offs exacerbated by the high noise floor. It ensures unwavering reliability—a paramount factor for mission success and safety—by intelligently selecting the best available path. This is essential, as it provides robust connectivity even when scarce power or geometric constraints make achieving the absolute highest reliability impractical.